# Identifying treatment response subgroups in observational time-to-event data

### v4

@@ 3.3 Training procedure
First, the model $W$ is trained to predict the binary treatment assignment probability by minimising the cross-entropy of receiving treatment. Then, training all other components relies on maximising the weighted log-likelihood introduced in Equation (2). The use of monotonic neural networks results in the efficient and exact computation of the log-likelihood as a forward path provides the estimated survival and automatic differentiation of the monotonic neural networks’ outcomes readily provides the derivative necessary for computing the log-likelihood (Jeanselme et al., 2022, 2023; Rindt et al., 2022).
@@ 4 Experimental analysis
As counterfactuals are unknown in observational data, we adopt—as is common practice in this research field—a synthetic dataset in which underlying survival distributions and group structure are known. As a real-world case study, Appendix C accompanies these synthetic results with the analysis of heterogeneity in adjuvant radiotherapy responses for patients diagnosed with breast cancer. Our code to produce the synthetic dataset, the model, and all experimental results is available on Github [^4] https://github.com/Jeanselme/CausalNeuralSurvivalClustering.
@@ 4.1 Data generation
We generate a population of $3,000$ equally divided into $K=3$ subgroups. We draw 10 covariates from normal distributions with different centres and survival distributions for each treatment regime following Gompertz distributions (Pollard & Valkovics, 1992) parametrised by group membership and individual covariates. Note that this setting relaxes the independence of outcomes and covariates assumed by the proposed model (Assumption 2). This choice reflects a setting in which survival distributions under treatment or control regimes are different functions of the covariates. This simulation is more likely to capture the complexity of real-world responses, in contrast to traditional evaluations of subgroup analysis, which often assume a linear treatment response. Further, we implement two treatment assignment scenarios: a randomised assignment in which treatment is independent of patient covariates, similar to RCTs (Randomised), and one in which treatment is a function of the patient covariates as in observational studies (Observational). Finally, non-informative censoring times are drawn following a Gompertz distribution. Details of the generative process of our synthetic dataset are deferred to Appendix B.1. Further, Appendix B.5 explores alternative datasets to demonstrate the flexibility of the proposed methodology.
@@ 4.2 Empirical settings
We compare the proposed approach (CSC) against the state-of-the-art Cox Mixtures with Heterogeneous Effect (CMHE (Nagpal et al., 2022a) [^5] Implemented in the Auton-Survival library (Nagpal et al., 2022b)) which uncovers treatment effect and baseline survival latent groups. This method uses an expectation maximisation framework in which each patient is assigned to a group for which a Cox model is then fitted. The central differences with our proposed approach are that CMHE (i) clusters patients in treatment effects subgroups and survival subgroups, and (ii) assumes a linear treatment response. This separation between survival and treatment response improves the model’s flexibility at the cost of interpretability as the number of groups grows exponentially, and subgroups of treatment effect are independent of survival. Further, the assumption of linear treatment response may hinder the discovery of subgroups with more complex responses. By contrast, our approach identifies treatment subgroups while considering survival without constraining the treatment response. We argue that these are key strengths of our method as a group not responding to treatment with low life expectancy would most benefit from alternative treatments, compared to a group with the same treatment responses but with better survival odds. Considering jointly non-linear treatment effects and survival, therefore, results in identifying more clinically relevant subgroups.
@@
For a fair comparison, we present three alternatives of CMHE: one with fixed $K=3$ survival subgroups, one with $L=3$ treatment effect subgroups, and one with $K=L=2$, which allows for a total of 4 subgroups. Crucially, these methods assume proportional hazards for each subgroup and do not account for the treatment non-randomisation.
@@
Additionally, we also compare our model against its unadjusted alternative (CSC Unadjusted), which uses the unweighted factual likelihood ($w_{i}=1$) and should suffer from non-randomisation of treatment assignment.
@@
Finally, we compare against two step-wise approaches in which clustering and treatment effect are modelled separately. First, we use an unsupervised clustering algorithm on the covariates, followed by a non-parametric estimate of the treatment effect as proposed in (Nagpal et al., 2022a), referred to as Kmeans + TE in the following. Specifically, we use a K-Means (Hartigan & Wong, 1979) to cluster the data, and we compute the difference between Kaplan-Meier (Kaplan & Meier, 1958) estimates between control and treated patients stratified by clusters. Second, we use a Virtual Twins approach to estimate individualised treatment effects and then cluster them. As proposed in the literature, we use a survival tree to estimate response under each treatment regime and then use a KMeans (MacQueen et al., 1967) on the estimated treatment effects, computed as the difference between survival estimates. Note that these methodologies assume treatment randomisation and the clustering is uninformed by the covariates. For details on training and hyperparameter optimisation, we refer to App. B.2.
@@
In the synthetic experiments, the subgroup structure is known. We measure the adjusted [^6] Random patient assignment results in an adjusted Rand-Index of 0. Rand-Index (Rand, 1971), which quantifies how the estimated assignment aligns with the known underlying group structure. Additionally, we use the integrated absolute error (IAE) between the treatment effect estimate and the ground truth, which measures how well we recover each subgroup’s treatment effect
[…]
@@ 4.3 Treatment effect recovery
Recovering the underlying number of subgroups.
For all methodologies, we must choose the number of subgroups a priori. An important question is, therefore, whether we can identify the true number of subgroups with our model. Figure 4 presents the average negative log-likelihood obtained by cross-validation for models with different numbers of subgroups $K$. The dotted lines represent the elbow heuristic (Thorndike, 1953), which identifies a change point in the explained variability of a clustering strategy, here considering the log-likelihood. Using this heuristic, the optimal choice for $K$ is $3$, which aligns with the underlying generative process. App. B.3 shows that when the number of clusters is misspecified, the estimated treatment effects are unstable, the variance associated with estimated groups is hence an additional tool to select the number of clusters. This data-driven choice of the number of subgroups $K$ is a crucial strength of our method, compared to classical two-stage analyses, which separate clustering from treatment effect estimates. In these methods, survival outcomes cannot directly guide the choice of $K$, whereas it is directly reflected in the proposed methodology’s performance.
@@
Next, we present the quantitative results of our method against the benchmark methods. Table 1 summarises the performance of the different methodologies under the two studied scenarios. Recall that $K$ denotes the number of treatment response subgroups, while in CMHE, the additional parameter $L$ describes the number of survival clusters.
@@
Model Rand-Index $\text{IAE}_{k}(t_{max})$ $k=1$ $k=2$ $k=3$ Randomised CSC 0.807 (0.055) 0.037 (0.008) 0.023 (0.008) 0.015 (0.003) CSC Unadjusted 0.802 (0.050) 0.039 (0.011) 0.021 (0.009) 0.014 (0.003) CMHE ($K=3$) 0.392 (0.034) 0.164 (0.009) 0.077 (0.004) 0.070 (0.005) CMHE ($L=3$) 0.255 (0.159) - 0.084 (0.005) - CMHE ($K=L$) 0.345 (0.104) 0.243 (0.042) - - KMeans + TE 0.000 (0.002) 0.210 (0.011) 0.091 (0.007) 0.142 (0.016) Virtual Twins 0.578 (0.081) 0.022 (0.006) 0.032 (0.011) 0.025 (0.010) Observational CSC 0.797 (0.043) 0.042 (0.011) 0.051 (0.026) 0.022 (0.004) CSC Unadjusted 0.742 (0.073) 0.047 (0.019) 0.030 (0.029) 0.020 (0.003) CMHE ($K=3$) 0.385 (0.022) 0.169 (0.012) 0.078 (0.005) 0.075 (0.008) CMHE ($L=3$) 0.190 (0.127) 0.192 (0.010) - 0.140 (0.009) CMHE ($K=L$) 0.454 (0.068) 0.210 (0.013) 0.095 (0.020) 0.188 (0.014) KMeans + TE 0.001 (0.004) 0.192 (0.016) 0.090 (0.019) 0.147 (0.016) Virtual Twins 0.438 (0.139) 0.050 (0.039) 0.034 (0.033) 0.051 (0.033)
@@
Table 1: 5-fold cross-validated performance averaged (with standard deviation in parenthesis) under the Randomised and Observational treatment simulation. Best performance per column and simulation scenario are marked in bold, second best in italic. ’-’ describes when the methodology diverges. Our proposed CSC method and its unadjusted variant best recover the underlying treatment responses, with the adjusted approach presenting the best performance in the observational setting.
@@
CSC outperforms all CMHE alternatives, the current state-of-the-art method. CMHE’s parametrisation and assumptions explain this difference. Critically, CMHE assumes (i) proportional hazards and (ii) a linear impact of treatment on log hazards. Neither of these two assumptions is likely to hold in real-world settings as mimicked by our synthetic data. In contrast, our approach does not constrain the treatment effect due to its flexible modelling of the survival function under both treatment regimes. Increasing the number of subgroups, as shown in $K=L$, improves the performance of CMHE in terms of clustering quality (Rand-Index) and the recovery of the underlying treatment effect (lower IAE), but still is inferior compared to our proposed methods. These experiments highlight the advantage of the proposed CSC method in uncovering subgroups of treatment responses due to the flexibility in modelling complex survival distributions under both treatment regimes without proportionality assumption.
@@
CSC presents the best performance in identifying the underlying subgroup. Our proposed CSC method outperforms all approaches on the Rand-Index. In particular, the two-step approaches do not identify the underlying subgroups when covariates and outcomes subgroups are unaligned. KMeans + TE identify subgroups that are independent of the outcome of interest, as shown by the Rand-Index. Consequently, the identified clusters differ in their treatment responses, as shown by the large IAE. The Virtual Twins approach presents a better capacity to identify the clusters of interest with better Rand-Index and IAE. However, the two-step approach hurts subgroup identification due to the disconnect between the clustering and input covariates. This disconnect leads to a significantly lower Rand-Index in comparison to CSC despite accurate treatment effect estimates.
@@
Treatment assignment correction improves subgroup identification in observational settings. The Observational simulation demonstrates the importance of correcting the likelihood under treatment non-randomisation. The two CSC alternatives present comparable performance in the randomised setting, as theoretically expected, due to a constant $w_{i}$ in this context. However, CSC better recovers the different groups as shown by the Rand-Index. Finally, Appendices B.5.1 to B.5.4 further validate the robustness of our method in settings with unequally distributed cluster sizes, population size, increased number of underlying clusters, different treatment rates and covariate structures.
Real-world heterogeneity analysis in adjuvant radiotherapy responses (App. C).
Using the Surveillance, Epidemiology, and End Results program(Seer) [^7] Available at https://seer.cancer.gov/, App. C investigates a real-world setting: we analyse which subgroups of breast cancer patients most benefit from adjuvant radiotherapy, an active area of research. Our analysis highlights a subgroup characterised by a higher number of distant lymph nodes that benefit from treatment. Despite the inherent limitations of the dataset, our exploratory method identifies a subgroup that alternative methods miss, potentially informing future clinical trials to validate or challenge this hypothesis. Critically, this demonstrates that our work offers a practical methodology for exploratory subgroup analysis of medical data to identify treatment effect heterogeneity in observational studies.
@@ 5 Conclusion
Understanding heterogeneity in treatment effect is essential for guiding clinical practice and informing the development of new treatments. Motivated principally by how real-world clinical guidelines depend on patient subgroups, this work introduces a neural network-based framework for identifying treatment-response subgroups in observational time-to-event data. By leveraging large-scale observational datasets and survival analysis with causal inference methods, our approach addresses a key gap in the literature: it moves beyond average or individualised treatment-effect estimation to reveal subgroups of treatment response.
Our empirical results on both synthetic and real-world datasets demonstrate that the proposed method successfully characterises latent subgroups with divergent treatment effects. While we acknowledge that our model depends on empirically unverifiable assumptions as often in causal inference, our focus on subgroup detection for hypothesis generation lessens this concern. Indeed, such subgroup identification serves as a starting point for more targeted RCTs, where further validation can (i) confirm the heterogeneity in responses, (ii) explore alternative treatment strategies in less responsive subgroups, and (iii) shape evidence-based clinical guidelines.
In doing so, this work contributes a flexible and practical tool for the design of future RCTs. By informing both the selection of patient populations most likely to benefit and avoiding treatments with limited or harmful effects, our framework offers a meaningful hypothesis-generating tool for clinicians, policymakers, and researchers striving to optimise patient care and resource allocation.

### v5

@@ 2.3 Training procedure
The model $W$ is trained to predict the binary treatment assignment probability by minimising the cross-entropy of receiving treatment. Then, all other components are trained by maximising the weighted log-likelihood introduced in Equation (2). The use of monotonic neural networks results in the efficient and exact computation of the log-likelihood, as a forward path estimates survival and automatic differentiation provides the derivative necessary for computing the log-likelihood (Jeanselme et al., 2022, 2023; Rindt et al., 2022).
@@ 3 Experimental analysis
As counterfactuals are unknown in observational data, we adopt the common practice of a synthetic dataset in which the underlying survival distributions and group structure are known. As a real-world case study, App. 4 accompanies these synthetic results with the analysis of heterogeneity in adjuvant radiotherapy responses for patients diagnosed with breast cancer. Our code to produce the synthetic dataset, the model, and all experimental results is available on Github [^5] https://github.com/Jeanselme/CausalNeuralSurvivalClustering.
@@ 3.1 Data generation
We generate a population of $3,000$ equally divided into $K=3$ subgroups. We draw 10 covariates from normal distributions with different means and survival distributions for each treatment regime following Gompertz distributions (Pollard and Valkovics, 1992) parametrised by group membership and individual covariates. Note that this setting breaks Asmp. 2 as survival distributions under treatment or control regimes are different functions of the covariates. This simulation is more likely to capture the complexity of real-world responses, in contrast to traditional evaluations of subgroup analysis, which often assume a linear treatment response. Further, we implement two treatment assignment scenarios: a randomised assignment in which treatment is independent of patient covariates, similar to RCTs (Randomised), and one in which treatment is a function of the patient covariates as in observational studies (Observational). Finally, non-informative censoring times are drawn following a Gompertz distribution. Details of the generative process for our synthetic dataset are deferred to App. C.1. Further, App. C.5 explores alternative datasets to demonstrate the robustness of the proposed methodology.
@@ 3.2 Empirical settings
We compare the proposed approach (CSC) against the state-of-the-art Cox Mixtures with Heterogeneous Effect (CMHE Nagpal et al. (2022a) [^6] Implemented in the Auton-Survival library (Nagpal et al., 2022b)), which uncovers treatment effect and baseline survival latent groups. This method uses an expectation maximisation framework in which each patient is assigned to a group for which a Cox model is then fitted. The central differences with our proposed approach are that CMHE (i) clusters patients in treatment effects subgroups and survival subgroups, and (ii) assumes a linear treatment response. This separation between survival ($M$) and treatment response ($L$) enhances the model’s flexibility at the expense of interpretability, as the total number of groups grows exponentially ($L\times M$), and subgroups of treatment effect become independent of survival. Further, the assumption of linear treatment response may hinder the discovery of subgroups with more complex responses. By contrast, our approach identifies treatment subgroups while considering survival without constraining the treatment response. We argue that these are key strengths of our method, as a group not responding to treatment with low life expectancy would most benefit from alternative treatments, compared to a group with the same treatment responses but with better survival odds. Considering jointly non-linear treatment effects and survival, therefore, results in identifying more clinically relevant subgroups.
@@
For a fair comparison, we present three alternatives of CMHE: one with fixed $M=3$ survival subgroups and $L=1$ treatment subgroups, one with $L=3$ treatment and $M=1$ survival subgroups, and one with $M=L=2$, which allows for a total of 4 subgroups. We consider these alternatives to obtain a total number of clusters close to the underlying number of subgroups $K$. Crucially, CMHE assumes proportional hazards for each subgroup and does not account for the treatment non-randomisation.
@@
As an ablation, we compare our model against its unadjusted alternative (CSC Unadjusted), which uses the unweighted factual likelihood ($w_{i}=1$) and therefore is not robust to non-random treatment assignment.
@@
Finally, we compare against two step-wise approaches in which clustering and treatment effect are trained separately. First, we use an unsupervised clustering algorithm on the covariates, followed by a non-parametric estimate of the treatment effect as proposed in Nagpal et al. (2022a), referred to as KMeans + TE in the following. Specifically, we use a KMeans (Hartigan and Wong, 1979) to cluster the data, and we compute the difference between Kaplan-Meier (Kaplan and Meier, 1958) estimates between control and treated patients stratified by clusters. Second, we use a Virtual Twins approach to estimate individualised treatment effects and then cluster them. As proposed in the literature, we use a survival tree to estimate response under each treatment regime and then use a KMeans (MacQueen et al., 1967) on the estimated treatment effects, computed as the difference between survival estimates. For details on training and hyperparameter optimisation, we refer to App. C.2.
@@
In the synthetic experiments, the subgroup structure is known. We measure the adjusted [^7] Random patient assignment results in an adjusted Rand-Index of 0. Rand-Index (Rand, 1971), which quantifies how the estimated subgroup assignment aligns with the generative group structure. Additionally, we use the integrated absolute error (IAE) between the treatment effect estimate and the ground truth, which measures how well the methods recover each subgroup’s treatment effect
$\text{IAE}_{k}(t)=\int_{0}^{t}\left|\hat{\tau}_{\hat{k}}(s)-\tau_{k}(s)\right|ds,$
@@ 3.3 Treatment effect recovery
Recovering the underlying number of subgroups.
For all methodologies, we must choose the number of subgroups $K$ a priori. An important question is, therefore, whether we can identify the underlying number of subgroups in a principled way using the proposed CSC. Figure 4 presents the average negative log-likelihood obtained by cross-validation for models with different numbers of subgroups $K$.
@@
The dotted lines represent the elbow heuristic (Thorndike, 1953), which identifies a change point in the explained variability, here considering the log-likelihood. Using this heuristic, the optimal choice for $K$ is $3$, which aligns with the generative process. App. C.3 shows that when the number of clusters is misspecified, the estimated treatment effects are unstable; the variance associated with estimated treatment effects is hence an additional tool for informing the choice of $K$. These data-driven approaches to select the number of subgroups $K$ are a crucial strength of our method, compared to classical two-stage analyses, which separate clustering from treatment effect estimates. In these methods, survival outcomes cannot directly guide the choice of $K$, whereas it is directly reflected in CSC’s performance.
@@
Table 1 summarises the performance of the different methodologies under the two studied scenarios. Recall that $L$ denotes the number of treatment response subgroups, while in CMHE, the additional parameter $M$ describes the number of survival clusters.
@@
Model Rand-Index $\text{IAE}_{k}(t_{max})$ $k=1$ $k=2$ $k=3$ Randomised CSC 0.807 (0.055) 0.037 (0.008) 0.023 (0.008) 0.015 (0.003) CSC Unadjusted 0.802 (0.050) 0.039 (0.011) 0.021 (0.009) 0.014 (0.003) CMHE ($L=3$) 0.392 (0.034) 0.164 (0.009) 0.077 (0.004) 0.070 (0.005) CMHE ($M=3$) 0.255 (0.159) - 0.084 (0.005) - CMHE ($M=L$) 0.345 (0.104) 0.243 (0.042) - - KMeans + TE 0.000 (0.002) 0.210 (0.011) 0.091 (0.007) 0.142 (0.016) Virtual Twins 0.578 (0.081) 0.022 (0.006) 0.032 (0.011) 0.025 (0.010) Observational CSC 0.797 (0.043) 0.042 (0.011) 0.051 (0.026) 0.022 (0.004) CSC Unadjusted 0.742 (0.073) 0.047 (0.019) 0.030 (0.029) 0.020 (0.003) CMHE ($L=3$) 0.385 (0.022) 0.169 (0.012) 0.078 (0.005) 0.075 (0.008) CMHE ($M=3$) 0.190 (0.127) 0.192 (0.010) - 0.140 (0.009) CMHE ($M=L$) 0.454 (0.068) 0.210 (0.013) 0.095 (0.020) 0.188 (0.014) KMeans + TE 0.001 (0.004) 0.192 (0.016) 0.090 (0.019) 0.147 (0.016) Virtual Twins 0.438 (0.139) 0.050 (0.039) 0.034 (0.033) 0.051 (0.033)
@@
Table 1: 5-fold cross-validated performance averaged (with standard deviation in parentheses) under the Randomised and Observational treatment simulation. Best performance per column and simulation scenario is marked in bold, second best in italic. ’-’ describes when the methodology diverges. Our proposed CSC method best recovers the underlying treatment responses in the observational setting.
@@
CSC outperforms all CMHE alternatives, the current state-of-the-art method. CMHE’s parametrisation and assumptions explain this difference. Critically, CMHE assumes (i) proportional hazards and (ii) a linear impact of treatment on log hazards. Neither of these two assumptions is likely to hold in real-world settings as mimicked by our synthetic data. In contrast, our approach does not constrain the treatment effect due to its flexible modelling of the survival function under both treatment regimes. Increasing the number of subgroups, as shown in $M=L$, improves the performance of CMHE in terms of clustering quality (Rand-Index) and the recovery of the underlying treatment effect (lower IAE), but still is inferior compared to our proposed methods. These experiments highlight the advantage of the proposed CSC method in uncovering subgroups of treatment responses due to the flexibility in modelling complex survival distributions under both treatment regimes.
@@
CSC presents the best performance in identifying the underlying subgroup. Our proposed CSC method outperforms all approaches on the Rand-Index. In particular, the two-step approaches do not identify the underlying subgroups when covariates and outcomes subgroups are unaligned. KMeans+TE identifies subgroups that are independent of treatment subgroups, as shown by the Rand-Index and the large IAE. The Virtual Twins approach presents a better capacity to identify the clusters of interest with better Rand-Index and IAE. However, the two-step approach hurts subgroup identification due to the disconnect between the clustering and input covariates. This disconnect leads to a significantly lower Rand-Index in comparison to CSC despite accurate treatment effect estimates.
@@
Treatment assignment correction improves subgroup identification in observational settings. The Observational simulation demonstrates the importance of correcting the likelihood under treatment non-randomisation. The two CSC alternatives present comparable performance in the randomised setting, as theoretically expected, due to a constant $w_{i}$ in this context. However, CSC better recovers the different groups as shown by the Rand-Index in the Observational setting.
4 Case study: Analysis of adjuvant radiotherapy responses
To illustrate how practitioners can use the proposed methodology to identify subgroups of treatment response in real-world observational data, we investigate how breast cancer patients respond differently to adjuvant radiotherapy using data from the Surveillance, Epidemiology, and End Results program (Seer). Following Lee et al. (2018); Danks and Yau (2022); Jeanselme et al. (2023), our analysis focuses on women who died from the condition or from cardiovascular diseases. To study the impact of adjuvant radiotherapy on survival outcomes after chemotherapy, we subselect patients with recorded chemotherapy. These criteria led to the selection of 239,855 women with 22 covariates measured at diagnosis, such as diagnosis year, grades, ethnicity, laterality, tumour size, and type (see Danks and Yau (2022) for further description).
Step 1: Select $K$.
From this population, our aim is to identify heterogeneous responses to adjuvant treatment. The first challenge is selecting the number of groups to use ($K$). We advise following medical actionability and considering the change in treatment effects and size of the subgroups when increasing this parameter. In the absence of experts’ insights, one may rely on the previously described elbow rule heuristic over $l_{F}$. Using Figure 7 in the Appendix, the negative log-likelihood presents an elbow for $K=2$.
Step 2: Identify subgroups.
Once $K$ is selected, one may use CSC to identify groups of patients who may benefit from adjuvant radiotherapy. This problem is central to patients’ management, as no evidence-based guidelines for adjuvant therapy exist Lazzari et al. (2023), making this setting more likely to meet the positivity assumption (Asmp. 4), which is necessary to study causality in observational data.
Figure 5 presents the identified treatment effect subgroups, and Table 2 summarises the mean covariates across the identified subgroups and the life expectancy gain when using adjuvant radiation, measured through the Restricted Mean Survival Time (RMST) (Royston and Parmar, 2013). The proposed methodology identifies two groups: one with limited treatment response and one characterised by larger HER2 prevalence and higher distant lymph node count, with a positive treatment response, gaining more than half a year of life expectancy over the five years following diagnosis. See App. D for the analysis of the treatment subgroups identified using the other baselines.
Step 3: Validate.
The proposed analysis identifies a group that benefits from adjuvant radiotherapy. However, our methodology remains a hypothesis-generating tool. An important next step would be to develop clinical trials to validate the identified subgroups, particularly due to potential confounding factors such as hormonal therapy (not available in this dataset), the temporal nature of treatment, and the plurality of treatment options. The quality of available covariates limits this analysis and serves only as an example for medical practitioners to identify subgroups of treatment responses from observational data.
@@ 5 Conclusion
Understanding heterogeneity in treatment effect is essential for guiding clinical practice and informing the development of new treatments. Motivated primarily by the reliance of clinical guidelines on patient subgroups, this work introduces a neural network-based framework for identifying treatment-response subgroups in observational time-to-event data. By leveraging large-scale observational datasets and survival analysis with causal inference methods, our approach addresses a key gap in the literature: it moves beyond average or individualised treatment-effect estimation to reveal subgroups of treatment response.
Our empirical results on both synthetic and real-world datasets demonstrate that the proposed method successfully characterises latent subgroups with divergent treatment effects. While we acknowledge that our model relies on empirically unverifiable assumptions, as is often the case in causal inference, our focus on subgroup detection for hypothesis generation mitigates this concern. Indeed, such subgroup identification serves as a starting point for more targeted RCTs, where further validation can (i) confirm the heterogeneity in responses, (ii) explore alternative treatment strategies in less responsive subgroups, and (iii) shape evidence-based clinical guidelines.
In doing so, this work provides a flexible and practical tool for designing future RCTs. By informing both the selection of patient populations most likely to benefit and avoiding treatments with limited or harmful effects, our framework offers a meaningful hypothesis-generating tool for clinicians, policymakers, and researchers striving to optimise patient care and resource allocation.